OpenStudy (anonymous):

PLEASE PLEASE HELP! A circle has a radius of 8 inches. Find the area of a sector of the circle if the sector has an arc that measures 45°. 2 sq. in.? 8 sq. in.? 16 sq. in.?

OpenStudy (anonymous):

Its 8 nches right?

OpenStudy (yttrium):

Do you know the formula in finding the are of a sector?

OpenStudy (yttrium):

I mean area

OpenStudy (nubeer):

You should start by converting the angle 45 into radian.. then formula for area is = 0.5*r^2* theta.. theta is the angle here.. you have the radius too given in the question.. so just plug in the values.

OpenStudy (yttrium):

The formula for finding the length of the sector:\[s = r theta\] For the area of a sector: \[A = (1/2) r ^{2} \theta\]

OpenStudy (anonymous):

We work with dragee !

OpenStudy (yttrium):

Well, nubeer is right. It must be converted first into radian.

OpenStudy (nubeer):

nop @E.ali .. we do these kind of problems in radian.. these formula can be applied when angle is in radian.

OpenStudy (anonymous):

We can work with cm and dragee ! Ha ?!:)

OpenStudy (anonymous):

@ nubeer ,8 in = cm ?

OpenStudy (nubeer):

doesn't matter if it's in cm or in.. it would work for both.. we just have to see for the angle.. if it's in radian or degree.

OpenStudy (anonymous):

Thank you guys

OpenStudy (nubeer):

but if u wanna know.. 1 inch = 2.5 cm

OpenStudy (anonymous):

If we work with Cm shold work with Dragee to and if work with in should work with Radian too !

OpenStudy (anonymous):

first of all, let's just make sure we understand the terminology, a sector is basically like a slice of the circle cut from the center out, like a slice of pie and the measure of the arc is the same as the measure of the central angle

OpenStudy (anonymous):